# Pretrained deep models outperform GBDTs in Learning-To-Rank under label scarcity

### v4

@@ 3.1 Evaluation of unsupervised pretraining techniques
In this subsection, we compare the pretraining strategies we discussed above in the following tabular LTR setting.
@@
Dataset. We use MSLRWEB30K (or MSLR for brevity) (Qin & Liu, 2013), Yahoo Set1 (or Set1 for brevity) (Chapelle & Chang, 2011), and Istella_S (Lucchese et al., 2016) (a smaller version of Istella (Dato et al., 2016)). In the tabular LTR literature, these three datasets are generally considered the standard; most recent papers evaluate (only) on these datasets (Qin et al., 2021; Ai et al., 2019; Pang et al., 2020; Yang et al., 2022; Yu, 2020). To simulate scarcity of labels, we let 0.1% of the training query groups in each dataset keep their labels, while the rest of the query groups have no labels. Dataset statistics are provided in Section A.2.1.
@@
Methodology. For the neural pretraining approaches (which include SimCLR-Rank, SimSiam, SimCLR, SubTab, VIME-self, DACL+, and SCARF), the base of the model is the tabular ResNet [^1] We emphasize that we do not prescribe any specific deep learning architecture choice; the choice for the best DL model may depend on the specific LTR task. Our proposed workflow of pretraining deep models in LTR is agnostic to the specific choice of DL architecture. Prior work has proposed DL architectures designed for click-through rate (CTR)/LTR based on factorization machines (Rendle, 2010), such as DeepFM and DCNv2 (Guo et al., 2017; Wang et al., 2021), which we compare to in Section 4.1.2 and Section 4.1.2 against our pretrained models. Note that it is unclear currently how to do unsupervised pretraining using these CTR/LTR specific architectures at the moment. (check the footnote for our explanation on the architecture choice). We also evaluate “SimCLR-Sample”, a neural pretraining baseline we create to demonstrate the value of negatives from the same query group. SimCLR-Sample is like SimCLR-Rank, but at the beginning of each epoch the items are first randomly permuted and grouped into fake query groups. Then we apply SimCLR-Rank. Thus SimCLR-Sample has the same time/space complexity as SimCLR-Rank. For the neural pretraining approaches (which include SimCLR-Rank, SimSiam, SimCLR, SubTab, VIME-self, DACL+, and SCARF), the base of the model is the tabular ResNet [^2] We emphasize that we do not prescribe any specific deep learning architecture choice; the choice for the best DL model may depend on the specific LTR task. Our proposed workflow of pretraining deep models in LTR is agnostic to the specific choice of DL architecture. Prior work has proposed DL architectures designed for click-through rate (CTR)/LTR based on factorization machines (Rendle, 2010), such as DeepFM and DCNv2 (Guo et al., 2017; Wang et al., 2021), which we compare to in Section 4.1.2 and Section 4.1.2 against our pretrained models. Note that it is unclear currently how to do unsupervised pretraining using these CTR/LTR specific architectures at the moment. (check the footnote for our explanation on the architecture choice). We pretrain on all the available query groups and finetune/train on only the labeled query groups.
@@
For the GBDT-based approaches, we use the implementation in lightgbm Ke et al. (2017) and adopt a tuning strategy similar to Qin et al. (2021). We report “Supervised GBDT” as GBDT utilizing only the labeled data, and in “Semi-supervised GBDT” we report the best among pseudo-labeled (Pseudo-Label, 2013) and PCA-enriched (Duh & Kirchhoff, 2008) GBDTs as well as the combination of the two techniques. We describe our experimental setup in detail in Section A.2.3. This includes hyperparameters, procedures for data augmentation, GBDT pseudo-labeling/PCA details, model training, and finetuning. For comparison to prior baselines on tabular data, our experimental setup is chosen to mimic the original setup in them.
@@
Results. Table 1 charts the test NDCG for each of the pretraining methods we study. SimCLR-Rank and SimSiam perform the best or the second best among these methods for most of the datasets. In particular, SimCLR-Rank outperforms SimCLR-Sample, demonstrating the value and efficiency of negatives from the same query group. As expected, SimCLR-Rank is also 7-14x faster than SimCLR/SCARF/DACL+ (Table 6 in Section A.3.1). We also find that the semi-supervised GBDT baselines do not perform well—surprisingly, they tend to achieve lower NDCG than the supervised GBDT baseline without semi-supervised learning. Given these advantages in terms of test NDCG and speed, we select SimCLR-Rank and SimSiam as the best overall pretraining methods for tabular LTR, and use them for our continued evaluation.
@@ 4 Empirical evaluation
Next we compare the best DL pretraining methods for tabular LTR (SimCLR-Rank and SimSiam) against GBDT and non-pretrained DL models on (1) the public datasets under different label scarcity patterns to simulate real-world scenarios [^3] Code is provided at https://anonymous.4open.science/r/ltr-pretrain-3C01/README.md., and (2) a private large-scale ranking dataset which is naturally label-scarce.
@@ 4.1 Public datasets results
For the public datasets (MSLRWEB30K, Yahoo Set1, Istella_S; details given in Section 3.1), we simulate label scarcity occuring in two scenarios: (1) crowdsourcing of labels for a limited number of query groups, and (2) using scarce implicit binary user feedback (e.g. clicks) as labels.
@@ 4.1.1 Simulated crowdsourcing of labels
Here we simulate a very practical setting where modelers use crowdsourcing to collect labels for only a fraction of the query groups due to limited resources.
@@
Methodology. We repeat the methodology from Section 3.1 for the GBDT and the SimCLR-Rank and SimSiam pretraining methods. For each dataset and labeling fraction, we report the test metric of the best GBDT ranker amongst semi-supervised GBDTs and supervised GBDTs under “GBDT”. Similarly, under “no pretrain DL” we report the best supervised DL model amongst tabular ResNet (Gorishniy et al., 2021), DeepFM (Guo et al., 2017), and DCNv2 (Wang et al., 2021) under the name “no pretrain DL”. Finally, we report the best pretrained ranker amongst SimCLR-Rank and SimSiam under the name “pretrained DL”. Pretrained DL rankers use tabular ResNet (see Section A.2.3 for more details). During hyperparameter tuning, we tune for NDCG and report the resulting test NDCG and test Outlier-NDCG.
Results. We compare pretrained rankers, non-pretrained DL rankers, and GBDTs on public datasets in Figure 1. The results for Outlier-NDCG are provided in Figure 8 in Section A.3.4. We find that pretrained rankers outperform non-pretrained methods (including GBDTs) on NDCG and Outlier-NDCG across all public datasets, up to a dataset-dependent fraction of QGs labeled, as shown in Figure 1 (for NDCG) and Figure 8 (for Outlier-NDCG). We also note that GBDTs have better NDCG than non-pretrained DL models in most of these regimes. This shows that there exist scenarios with a limited supply of labeled query groups, where self-supervised pretraining is the factor that allows deep learning models to outperform GBDTs. We also provide a detailed a comparison between various semi-supervised and supervised GBDT models in Section A.4
@@ 4.1.2 Simulated implicit binary user feedback
Dataset. To simulate this scenario, we follow the methodology from Yang et al. (2022), to generate stochastic binary labels from the true relevance labels for training and validation sets of each of the public datasets. Note that we still use the true (latent) relevance labels in the test set for evaluation. This models a scenario where we observe binary labels that are noisy observations of a true latent label, but the task is to use these noisy labels to learn to rank according to the true labels. The details of this binarizing transformation are provided in Section A.2.4. In this transformation, a parameter, $\tau_{\text{target}}$, implicitly controls how sparse the binary labels are (larger is more sparse). We select $\tau_{\text{target}}=4.5$, to obtain 8.9%, 5.3%, and 25.2% labeled query groups in MSLR, Yahoo Set1, Istella_S datasets, respectively. We present detailed data statistics in Table 11 of Appendix.
Table 2: Main result (Simulated implicit binary user feedback): When $X$(MSLRWEB30K=8.9%, Yahoo Set1=5.3%, and in Istella_S=25.2%) of query groups are assigned a few binary labels, pretrained DL method outperform both GBDT and non-pretrained DL models. We generate labels with $\tau_{\text{target}}=4.5$ (Section 4.1.2). means that pretrained models are significantly better in NDCG than non-pretrained models, measured by a t-test with significance $p<0.05$.
@@
Methodology. We reuse the methodology of Section 4.1.1, except we use a linear scoring head for pretrained models to improve stability and performance (see Section A.2.3). Similar to Section 4.1.1, we report the best of multiple models for each method and dataset (see Section 4.1.1 for details).
Results. In Table 2, we present the test NDCG and Outlier-NDCG metrics of the models under this setting. We see that pretrained rankers significantly outperform GBDTs and non-pretrained DL rankers in this simulated binary user feedback setting with up to, even with 25.2% labeled query groups (Istella_S). Notice that for MSLRWEB30K and Yahoo Set1 datasets, the second best model is the GBDT, and the non-pretrained DL ranker has a much worse Outlier-NDCG. This aligns with the recent observation that GBDT models perform better than DL models in supervised training with datasets containing irregularities McElfresh et al. (2023). However, here pretraining helps DL rankers close this gap and at times beat GBDTs in outlier performance. We also present results for $\tau_{\text{target}}\in\{4.25,5.1\}$ in Section A.3.5. There, we find that making the labels sparser ($\tau_{\text{target}}=5.1$) increases the relative improvement of pretrained DL rankers, while in a less sparse setting ($\tau_{\text{target}}=4.25$) GBDTs dominate. This strongly suggests that pretrained DL models outperform GBDTs when learning from relatively rare implicit binary user feedback, which is common in search and recommendation systems.
@@ 4.2 Large-scale real-world dataset results
In this section we test whether the results from the simulated label scarce settings above, especially Section 4.1.2, translate to a large-scale real-world LTR problem.
Dataset. Our proprietary dataset is derived from the online shopping logs of a large online retailer, numbering in the tens of millions of query groups. In this dataset, query groups consist of items presented to a shopper if they enter a search query. We assign each item a purchase label: label is 1 if the shopper purchased the item, or 0 if the shopper did not purchase the item. Hence, this is a real-world instance of the implicit binary user feedback setting simulated in Section 4.1.2. Further, only about 10% of the query groups have items with non-zero labels. See Section A.2.1 for more dataset details.
Methodology. We compare the performance of pretrained DL models–SimCLR-Rank and SimSiam–with GBDT and non-pretrained DL models. SimSiam and SimCLR-Rank are first pretrained on all the query groups before being finetuned on the labeled query groups with some non-zero purchase labels. The GBDT and the non-pretrained models are only trained on the labeled data. More details about these models and training strategies is given in Section A.2.5.
Table 3: Main result (Real-world dataset): Results on a large-scale proprietary online shopping dataset. We give percentage improvement of rankers pretrained by SimSiam and SimCLR-Rank over GBDTs. Unsupervised pre-training improves (1) overall (full dataset) performance (internally, 2% is a big improvement), and (2) outlier performance.
@@
$\Delta$change w.r.t. GBDT
@@
Results. Our results are summarized in Table 3 and all numbers are given as relative percentage improvements ($\Delta\%$) over the GBDT ranker. First, we note that all the DL models outperform GBDT models in terms of NDCG. Amongst them, the pretrained model using SimSiam performs the best with a gain of over $5.5\%$. This is substantial, given that in large-scale industry datasets, even a 2% improvement in NDCG is considered significant. In terms of Outlier-NDCG, the story is very different. Here, we see that the non-pretrained DL model and SimCLR-Rank performs up to -7% worse than the GBDT. However, SimSiam pretraining allows the DL ranker to overcome this shortcoming and beat the outlier performance of the GBDT. This verifies our hypothesis from Section 4.1.2 that pretraining can help DL models outperform GBDTs on large-scale real-world LTR problems, especially with implicit user feedback. Lastly, we see a significant performance difference between SimSiam and SimCLR-Rank. We will investigate this using an ablation study with public datasets later in (Section 4.3).
@@ 4.3 Ablations: Choices for pretrained models in LTR
Figure 4: t-SNE plots of SimCLR-Rank, SimSiam, and fully supervised embeddings on Yahoo Set1 datasets. Relevance label is indicated by marker size/color. We see that SimCLR-Rank and SimSiam embeddings are qualitatively very different. Both SimCLR-Rank and SimSiam embeddings do not sort the data as well as fully supervised embeddings, which most effectively groups data points with relevance label 4 together.
Methodology. The dataset and methodology follows that of Section 4.1.1, except we let (1) SimCLR-Rank use the Gaussian augmentation with scale 1, and (2) SimSiam use the zeroing augmentation with probability 0.1. These augmentations generally perform well for their respective methods. Experiments illustrating the effect of each augmentation method are included in Section A.3.6.
Results. In Figure 3 we find that SimCLR-Rank performs better on MSLRWEB30K and Istella_S, while SimSiam performs much better on Yahoo Set1 when labels are scarce. Figure 4 shows the t-SNE projections of sub-sampled item embeddings learned by SimCLR-Rank, SimSiam, and pre-final layer of a fully supervised model for Yahoo Set1. We see that SimCLR-Rank and SimSiam also have qualitatively different embeddings: SimSiam’s embeddings are more clustered while SimCLR-Rank’s embeddings are less clustered. These observations raise the question of whether we can constructively combine them. We give similar t-SNE plots of other datasets in Section A.3.3
@@
Methodology. To produce a “SimCLR-Rank + SimSiam” model, we finetune a linear layer over the embeddings from a finetuned SimCLR-Rank concatenated with the embeddings from a finetuned SimCLR-Rank model. More details about how we combine the models is given in Section A.3.7.
Results. In Figure 3 we find that “SimCLR-Rank + SimSiam” performs equal to or better than SimCLR-Rank and SimSiam across many different percentages of labeled query groups. While the combined model does not uniformly dominate across all labeled query group percentages, it is a promising first step towards a single unified method for unsupervised pretraining in ranking.
@@
Methodology. We use the methodology from Section 3.1 to compare three finetuning strategies: linear probing (LP), full-finetuning (FF), and multilayer probing (MP) for SimSiam and SimCLR-Rank. MP tunes a 3-layer MLP head on top of the frozen pretrained embeddings. More experimental details are provided in Section A.2.3.
Results. Table 4 provides results on MSLRWEB30K dataset. We find that in terms of NDCG of both SimCLR-Rank and SimSiam, LP is the worst and FF is the best. Interestingly, MP comes as a close second in SimCLR-Rank, but it performs poorly on SimSiam. Our results on other datasets are similar (Section A.3.2). We thus recommend full fine-tuning as a stable and performant strategy.
To explain the above observations, we we can look again at the t-SNE projections of the embeddings generated by the SimSiam and SimCLR-Rank encoders and the pre-final layer of a fully supervised DL model trained a randomly sampled 1000 samples from the full training set (Figure 4). These plots offer the following explanations for the two phenomena. (1) The embeddings produced by the SimSiam/SimCLR-Rank are not sorted by relevance score in the projection space, unlike the fully supervised encoder. This suggests that a linear ranker cannot use the pretrained embeddings to predict the true relevance labels. (2) MP works better for SimCLR-Rank than SimSiam because SimCLR-Rank’s embeddings are more evenly spread out (less collapsed) than those of SimSiam. This allows an MLP trained on the SimCLR-Rank embeddings to distinguish the items with different relevance labels.
Table 4: Finetuning: Comparison of fine tuning strategies on MSLRWEB30K dataset (LP = Linear Probing, MP = Multilayer Probing, FF = Full Finetuning). Full finetuning consistently performs best or near-best, while linear probing and multilayer probing perform poorly.

### v5

@@ 3.1 Evaluation of unsupervised pretraining techniques
In this subsection, we compare the discussed pretraining strategies in the following tabular LTR setting.
@@
Dataset. We use MSLRWEB30K (or MSLR for brevity) (Qin & Liu, 2013), Yahoo Set1 (or Set1 for brevity) (Chapelle & Chang, 2011), and Istella_S (Lucchese et al., 2016) (a smaller version of Istella (Dato et al., 2016)). In the tabular LTR literature, these three datasets are generally considered the standard; most recent papers evaluate (only) on these datasets (Qin et al., 2021; Ai et al., 2019; Pang et al., 2020; Yang et al., 2022; Yu, 2020). To simulate scarcity of labels, we let 0.1% of the training query groups in each dataset keep their labels, while the rest of the query groups have no labels. Dataset statistics are provided in Section˜A.2.1.
@@
Methodology. For the neural pretraining approaches (which include SimCLR-Rank, SimSiam, SimCLR, SubTab, VIME-self, DACL+, SAINT, and SCARF), the base of the model is the tabular ResNet [^1] We emphasize that we do not prescribe any specific deep learning architecture choice; the choice for the best DL model may depend on the specific LTR task. Our proposed workflow of pretraining deep models in LTR is agnostic to the specific choice of DL architecture. Prior work has proposed DL architectures designed for click-through rate (CTR)/LTR based on factorization machines (Rendle, 2010), such as DeepFM and DCNv2 (Guo et al., 2017; Wang et al., 2021), which we compare to in Section 4.1.2 and Section 4.1.2 against our pretrained models. Note that it is unclear currently how to do unsupervised pretraining using these CTR/LTR specific architectures at the moment. (check the footnote for our explanation on the architecture choice). We also evaluate “SimCLR-Sample”, a neural pretraining baseline we create to demonstrate the value of negatives from the same query group. SimCLR-Sample is like SimCLR-Rank, but at the beginning of each epoch the items are first randomly permuted and grouped into fake query groups. Then we apply SimCLR-Rank. Thus SimCLR-Sample has the same time/space complexity as SimCLR-Rank. We pretrain on all the available query groups and finetune/train on only the labeled query groups.
@@
For the GBDT-based approaches, we use the implementation in lightgbm Ke et al. (2017) and adopt a tuning strategy similar to Qin et al. (2021). We report “Supervised GBDT” as GBDT utilizing only the labeled data, and in “Semi-supervised GBDT” we report the best among pseudo-labeled (Pseudo-Label, 2013) and PCA-enriched (Duh & Kirchhoff, 2008) GBDTs as well as the combination of the two techniques. We also evaluate GBDTs enriched with SimCLR-Rank and SimSiam embeddings as additional features, which we call “SimCLR-Rank + GBDT” and “SimSiam + GBDT”. We describe our experimental setup in detail in Section˜A.2.3. This includes hyperparameters, procedures for data augmentation, GBDT pseudo-labeling/PCA details, model training, and finetuning. For comparison to prior baselines on tabular data, our experimental setup is chosen to mimic the original setup in them.
@@
Results. Table˜1 charts the test NDCG for each of the pretraining methods we study. We highlight a few observations: (1) The pretraining-based DL methods almost all outperform GBDTs (both supervised and semi-supervised) in this label-scarce setting. Surprisingly, the semi-supervised GBDT baselines tend to achieve lower NDCG than the supervised GBDT baseline without semi-supervised learning. (2) When GBDTs are allowed to use representations learned from our pretraining methods, their performance consistently improves, though they do not outperform pretraining with DL. (3) SimCLR-Rank and SimSiam perform the best or the second best among these methods for most of the datasets. In particular, SimCLR-Rank outperforms SimCLR-Sample, demonstrating the value and efficiency of negatives from the same query group. As expected, SimCLR-Rank is also 7-14x faster than SimCLR/SCARF/DACL+ (Table˜6 in Section˜A.3.1). Given these advantages in terms of test NDCG and speed, we select SimCLR-Rank and SimSiam as the best overall pretraining methods for tabular LTR, and use them for our continued evaluation.
@@ 4 Empirical evaluation
Next we compare the best DL pretraining methods for tabular LTR (SimCLR-Rank and SimSiam) against GBDT and non-pretrained DL models on (1) the public datasets under different label scarcity patterns to simulate real-world scenarios [^2] Code is provided at https://github.com/houcharlie/ltr-pretrain/., and (2) a private large-scale ranking dataset which is naturally label-scarce.
@@ 4.1 Public datasets results
For the public datasets (MSLRWEB30K, Yahoo Set1, Istella_S; details given in Section˜3.1), we simulate label scarcity occurring in two scenarios: (1) crowdsourcing of labels for a limited number of query groups, and (2) using scarce implicit binary user feedback (e.g. clicks) as labels.
@@ 4.1.1 Simulated crowdsourcing of labels
We simulate a setting where labels are crowdsourced for only some query groups due to limited resources.
@@
Methodology. We repeat the methodology from Section˜3.1 for the GBDT and the SimCLR-Rank and SimSiam pretraining methods. For each dataset and labeling fraction, we report the test metric of the best GBDT ranker among semi-supervised GBDTs and supervised GBDTs under “GBDT”. Similarly, under “no pretrain DL” we report the best supervised DL model amongst tabular ResNet (Gorishniy et al., 2021), DeepFM (Guo et al., 2017), and DCNv2 (Wang et al., 2021) under the name “no pretrain DL”. Finally, we report the best pretrained ranker amongst SimCLR-Rank and SimSiam under the name “pretrained DL”. Pretrained DL rankers use tabular ResNet (see Section˜A.2.3 for more details). During hyperparameter tuning, we tune for NDCG and report the resulting test NDCG and test Outlier-NDCG.
Results. We compare pretrained rankers, non-pretrained DL rankers, and GBDTs on public datasets in Figure˜1. The results for Outlier-NDCG are provided in Figure˜8 in Section˜A.3.4. We find that pretrained rankers outperform non-pretrained methods (including GBDTs) on NDCG and Outlier-NDCG across all public datasets, up to a dataset-dependent fraction of QGs labeled, as shown in Figure˜1 (for NDCG) and Figure˜8 (for Outlier-NDCG). We also note that GBDTs have better NDCG than non-pretrained DL models in most of these regimes. This shows that there exist scenarios with a limited supply of labeled query groups, where self-supervised pretraining is the factor that allows deep learning models to outperform GBDTs. We provide a detailed comparison between semi-supervised and supervised GBDT models in Section˜A.4.
@@ 4.1.2 Simulated implicit binary user feedback
Dataset. To simulate this scenario, we follow the methodology from Yang et al. (2022), to generate independent stochastic binary labels for each item from its true relevance label for training and validation sets of each of the public datasets. Note that we still use the true (latent) relevance labels in the test set for evaluation. This models a scenario where we observe binary labels that are noisy observations of a true latent label, but the task is to use these noisy labels to learn to rank according to the true labels. The details of this binarizing transformation are provided in Section˜A.2.4. In this transformation, a parameter, $\tau_{\text{target}}$, implicitly controls how sparse the binary labels are (larger is more sparse). We select $\tau_{\text{target}}=4.5$, to obtain 8.9%, 5.3%, and 25.2% labeled query groups which contain at least one positive label in MSLR, Yahoo Set1, Istella_S datasets, respectively. We present detailed data statistics in Table˜11 of Appendix.
Table 2: Main result (Simulated implicit binary user feedback): When XX% (MSLRWEB30K=8.9%, Yahoo Set1=5.3%, and in Istella_S=25.2%) of query groups are assigned a few binary labels, pretrained DL method outperform both GBDT and non-pretrained DL models. We generate labels with $\tau_{\text{target}}=4.5$ (Section˜4.1.2). means that pretrained models are significantly better in NDCG than non-pretrained models, measured by a t-test with significance $p<0.05$.
@@
Methodology. We reuse the methodology of Section˜4.1.1, except we use a linear scoring head for pretrained models to improve stability and performance (see Section˜A.2.3). Similar to Section˜4.1.1, we report the best of multiple models for each method and dataset (see Section˜4.1.1 for details).
Results. In Table˜2, we present the test NDCG and Outlier-NDCG metrics of the models under this setting. We see that pretrained rankers significantly outperform GBDTs and non-pretrained DL rankers in this simulated binary user feedback setting with up to, even with 25.2% labeled query groups (Istella_S). For MSLRWEB30K and Yahoo Set1 datasets, the second best model is the GBDT, and the non-pretrained DL ranker has a much worse Outlier-NDCG. This aligns with recent observations that GBDTs perform better than DL models in supervised training with datasets containing irregularities McElfresh et al. (2023). However, pretraining helps DL rankers close this gap, and at times beat GBDTs in outlier performance. We also present results for $\tau_{\text{target}}\in\{4.25,5.1\}$ in Section˜A.3.5. There, we find that making the labels sparser ($\tau_{\text{target}}=5.1$) increases the relative improvement of pretrained DL rankers, while in a less sparse setting ($\tau_{\text{target}}=4.25$) GBDTs dominate. This suggests that pretrained DL models outperform GBDTs when learning from rare implicit binary user feedback, as is common in search and recommendation systems.
@@ 4.2 Large-scale real-world dataset results
In this section we test whether the results from the simulated label scarce settings above, especially Section˜4.1.2, translate to a large-scale real-world LTR problem.
Dataset. Our proprietary dataset is derived from the online shopping logs of a large online retailer, numbering in the tens of millions of query groups. In this dataset, query groups consist of items presented to a shopper if they enter a search query. We assign each item a purchase label: label is 1 if the shopper purchased the item, or 0 if the shopper did not purchase the item. Hence, this is a real-world instance of the implicit binary user feedback setting simulated in Section˜4.1.2. Further, only about 10% of the query groups have items with non-zero labels. See Section˜A.2.1 for more dataset details.
Methodology. We compare the performance of pretrained DL models–SimCLR-Rank and SimSiam–with GBDT and non-pretrained DL models. SimSiam and SimCLR-Rank are first pretrained on all the query groups before being finetuned on the labeled query groups with some non-zero purchase labels. The GBDT and the non-pretrained models are only trained on the labeled data. More details about these models and training strategies is given in Section˜A.2.5.
Table 3: Main result (Real-world dataset): $\%$ improvement of rankers pretrained by SimSiam and SimCLR-Rank over GBDTs on a large-scale proprietary online shopping dataset. Unsupervised pretraining improves performance (1) overall (full dataset) (internally, 2% is a big improvement), and (2) on outliers.
@@
$\Delta$% change w.r.t. GBDT
@@
Results. Our results are summarized in Table˜3 and all numbers are given as relative percentage improvements ($\Delta\%$) over the GBDT ranker. First, note that all the DL models outperform GBDT models in terms of NDCG. Amongst them, the pretrained model using SimSiam performs the best with a gain of over $5.5\%$. This is substantial, given that in large-scale industry datasets, even a 2% improvement in NDCG is considered significant. In terms of Outlier-NDCG, the story is very different. Here, the non-pretrained DL model and SimCLR-Rank performs up to -7% worse than the GBDT. However, SimSiam pretraining allows the DL ranker to overcome this shortcoming and beat the outlier performance of the GBDT. This verifies our hypothesis from Section˜4.1.2 that pretraining can help DL models outperform GBDTs on large-scale real-world LTR problems, especially with implicit user feedback. Lastly, we see a significant performance difference between SimSiam and SimCLR-Rank. We investigate this in an ablation with public datasets in Section˜4.3.
@@ 4.3 Ablations: Choices for pretrained models in LTR
Figure 4: t-SNE plots of SimCLR-Rank, SimSiam, and a non-pretrained model trained on 0.1% of the data for the Yahoo Set1 dataset. Both SimCLR-Rank and SimSiam use the zeros augmentation ($p=0.1$) to remove augmentation choice as a confounding variable. Relevance label is indicated by marker size/color. We see that SimCLR-Rank and SimSiam embeddings are qualitatively very different. Qualitatively, SimCLR-Rank and SimSiam embeddings separate different relevance labels better than the non-pretrained model. For this setting, SimSiam has the best downstream performance, and also gives the most clustered embeddings.
Methodology. The dataset and methodology follows that of Section˜4.1.1, except we let (1) SimCLR-Rank use the Gaussian augmentation with scale 1, and (2) SimSiam use the zeroing augmentation with probability 0.1. These augmentations generally perform well for their respective methods. Experiments illustrating the effect of each augmentation method are included in Section˜A.3.6.
Results. In Figure˜3, SimCLR-Rank performs better on MSLRWEB30K and Istella_S, while SimSiam performs better on Yahoo Set1 when labels are scarce. Figure˜4 shows the t-SNE projections of sub-sampled item embeddings learned by SimCLR-Rank, SimSiam, and pre-final layer of a fully supervised model for Yahoo Set1. SimSiam exhibit qualitatively more clustered embeddings than SimCLR-Rank, particularly for high relevance labels. Similar t-SNE plots for other datasets are in Section˜A.3.3.
@@
Methodology. To produce a “SimCLR-Rank + SimSiam” model, we finetune a linear layer over the embeddings from a finetuned SimCLR-Rank concatenated with the embeddings from a finetuned SimCLR-Rank model. More details about how we combine the models is given in Section˜A.3.7.
Results. In Figure˜3 we find that “SimCLR-Rank + SimSiam” performs equal to or better than SimCLR-Rank and SimSiam across many different percentages of labeled query groups. While the combined model does not uniformly dominate across all labeled query group percentages, it is a promising first step towards a single unified method for unsupervised pretraining in ranking.
@@
Methodology. We use the methodology from Section˜3.1 to compare 3 finetuning strategies: linear probing (LP), full-finetuning (FF), and multilayer probing (MP) for SimSiam and SimCLR-Rank. MP tunes a 3-layer MLP head on top of the frozen pretrained embeddings. Experimental details are in Section˜A.2.3.
Results. Table˜4 provides results on MSLRWEB30K dataset. We find that in terms of NDCG of both SimCLR-Rank and SimSiam, LP is the worst and FF is the best. Interestingly, MP comes as a close second in SimCLR-Rank, but it performs poorly on SimSiam. Our results on other datasets are similar (Section˜A.3.2). We thus recommend full fine-tuning as a stable and performant strategy. To explain the above observations, we we can look again at the t-SNE projections of the embeddings generated by the SimSiam and SimCLR-Rank encoders (pretrained using the zeros augmentation with $p=0.1$ to remove augmentation choice as a confounding factor, finetuned on 0.1% of the data) and the pre-final layer of a model trained only on 0.1% of the data on a randomly sampled 1000 samples from the full training set (Figure˜4).
These plots offer the following explanations for the two phenomena. (1) The embeddings produced by the SimSiam/SimCLR-Rank are not sorted by relevance score in the projection space, unlike the fully supervised encoder. This suggests that a linear ranker cannot use the pretrained embeddings to predict the true relevance labels. (2) MP works better for SimCLR-Rank than SimSiam because SimCLR-Rank’s embeddings are more evenly spread out (less collapsed) than those of SimSiam. This allows an MLP trained on the SimCLR-Rank embeddings to distinguish the items with different relevance labels.
Table 4: Finetuning: Comparison of fine tuning strategies on MSLRWEB30K dataset (LP = Linear Probing, MP = Multilayer Probing, $F$ = Full Finetuning). Full finetuning consistently performs best or near-best, while linear probing and multilayer probing perform poorly.